# Arithmetic Properties of Generalized Cubic and Overcubic Partitions

### v1

@@ 1. Introduction
(1.3) $a_{37}(43n+12)$ $\equiv 0\pmod{43},$
(1.4) $a_{41}(47n+21)$ $\equiv 0\pmod{47},$
(1.5) $a_{53}(59n+56)$ $\equiv 0\pmod{59},$
(1.6) $a_{61}(67n+19)$ $\equiv 0\pmod{67},$
(1.7) $a_{65}(71n+32)$ $\equiv 0\pmod{71},$
(1.8) $a_{73}(79n+62)$ $\equiv 0\pmod{79},$
(1.9) $a_{77}(83n+79)$ $\equiv 0\pmod{83}.$
@@
After we finished the proof of Theorem 1.1, Guadalupe [Gua25] updated his preprint on arXiv, which actually generalizes the above result. However, he uses elementary techniques to achieve this, so we keep our proof in this version.
@@
We also give a complete characterization of the $\bar{a}_{c}(n)$ function modulo $4$ and $8$ in the following results.
@@
Theorem 1.6.
For all $n\geq 1$, we have
$\bar{a}_{c}(n)\equiv\begin{cases}2\pmod{8}&\text{if }n=k^{2},\,k\in\mathbb{Z},\\ 6(c+1)\pmod{8}&\text{if }n=2(2k)^{2},\,k\in\mathbb{Z},\\ 2(c+1)\pmod{8}&\text{if }n=2(2k-1)^{2},\,k\in\mathbb{Z},\\ 2(c+1)(c+2)\pmod{8}&\text{if }n=4k^{2},\,k\in\mathbb{Z},\\ 4(c+1)\pmod{8}&\text{if }n=k^{2}+2\ell^{2},\,k,\ell\in\mathbb{Z},\\ 0\pmod{8}&\text{otherwise.}\end{cases}$
Remark 1.7.
Theorem 1.6 generalizes a result of Sellers [Sel14, Theorem 2.7].
Theorems 1.4 and 1.6 are proved in Section 4 using elementary techniques.
@@
For all $n\geq 0$ and $i\geq 1$, we have
(1.11) $\bar{a}_{2i}(4n+3)$ $\equiv 0\pmod{4},$
(1.12) $\bar{a}_{2i}(8n+6)$ $\equiv 0\pmod{4},$
(1.13) $\bar{a}_{2i}(8n+5)$ $\equiv 0\pmod{4},$
(1.14) $\bar{a}_{2i}(8n+7)$ $\equiv 0\pmod{8}.$
Theorem 1.9.
For all $n\geq 0$, $i\geq 1$, and $\alpha\geq 0$, we have
$\bar{a}_{2i}(2^{\alpha}(4n+3))\equiv 0\pmod{4}.$
Theorem 1.10.
@@
(1.15) $\bar{a}_{2^{k}i-2^{k-1}-2}(8n+5)$ $\equiv 0\pmod{2^{k+1}},$
(1.16) $\bar{a}_{2^{k}i-2^{k-1}-2}(8n+7)$ $\equiv 0\pmod{2^{k+2}}.$
@@
(1.17) $\bar{a}_{2^{k}i+\frac{2^{k}-6}{2}}(8n+7)$ $\equiv 0\pmod{2^{k+3}},$
(1.18) $\bar{a}_{2^{k}i+\frac{2^{k}-2}{2}}(8n+7)$ $\equiv 0\pmod{2^{k+3}}.$
@@
(1.19) $\bar{a}_{9i}(9n+6)$ $\equiv 0\pmod{12},$
(1.20) $\bar{a}_{9i+4}(9n+6)$ $\equiv 0\pmod{12},$
(1.21) $\bar{a}_{9i+8}(9n+6)$ $\equiv 0\pmod{12}.$
Theorems 1.8, 1.9, 1.10, 1.11, and 1.12 are proved in Section 5 using elementary techniques.
@@
which means that almost all the coefficients of $A$ are divisible by $M$. It turns out that the partition function $\bar{a}_{c}(n)$ also satisfies such results. Specifically, we prove the following results in this section.
@@
$\{n\in\mathbb{N}:\bar{a}_{2^{\alpha}m}\equiv 0\pmod{2^{k}}\}$
@@
$\lim_{X\to\infty}\frac{\#\{0\leq n\leq X:\bar{a}_{2^{\alpha}m}\equiv 0\pmod{2^{k}}\}}{X}=1.$
@@
$\{n\in\mathbb{N}:\bar{a}_{2^{\alpha}m}\equiv 0\pmod{3^{k}}\}$
@@
$\lim_{X\to\infty}\frac{\#\{0\leq n\leq X:\bar{a}_{2^{\alpha}m}\equiv 0\pmod{3^{k}}\}}{X}=1.$
Theorem 1.15.
Let $k\geq 1$, $a_{1},a_{2},\dots,a_{m}$ be positive integers for some positive integer $m$, for $t=p_{1}^{a_{1}}\cdots p_{m}^{a_{m}}$ where $p_{i}$’s are prime numbers greater than or equal to $5$ with $2p_{i}^{k+a_{i}}\geq 5t$, the set
$\{n\in\mathbb{N}:\bar{a}_{t}\equiv 0\pmod{p_{i}^{k}}\}$
has arithmetic density $1$, namely,
$\lim_{X\to\infty}\frac{\#\{0\leq n\leq X:\bar{a}_{t}\equiv 0\pmod{p_{i}^{k}}\}}{X}=1.$
@@
Theorems 1.13, 1.14, 1.15, and 1.16 are proved in Section 6.
@@ 2.1. Various results on modular forms
$\Gamma$ :=$:=\left\{\begin{bmatrix}a&b\\ c&d\end{bmatrix}:a,b,c,d\in\mathbb{Z},ad-bc=1\right\},$
$\Gamma_{\infty}$ :=$:=\left\{\begin{bmatrix}1&n\\ 0&1\end{bmatrix}\in\Gamma:n\in\mathbb{Z}\right\}.$
@@
[…]
@@
If $\chi$ is a Dirichlet character modulo $N$, then a form $f(x)\in M_{l}(\Gamma_{1}(N))$ has Nebentypus character $\chi$ if
$f\left(\frac{az+b}{cz+d}\right)=\chi(d)(cz+d)^{l}f(z),$
@@
Let $k$ be an integer and $g(z)=\sum_{n=0}^{\infty}a(n)q^{n}$ a modular form of weight $k$ for $\Gamma_{0}(N)$. For any given positive integer $m$, if $a(n)\equiv 0\pmod{m}$ holds for all $n\leq\dfrac{kN}{12}\prod\limits_{p\leavevmode\nobreak\ \text{prime},\leavevmode\nobreak\ p\,|\,N}\leavevmode\nobreak\ \left(1+\dfrac{1}{p}\right)\,$ then $a(n)\equiv 0\pmod{m}$ holds for any $n\geq 0$.
@@
(2.1) $f(z)\,|\,T_{p}:=\sum_{n=0}^{\infty}\left(a(pn)+\chi(p)p^{\ell-1}a\left(\frac{n}{p}\right)\right)q^{n}.$
@@
#$\#\left\{0<n\leq X:b(n)\not\equiv 0\pmod{r}\right\}=\mathcal{O}\left(\frac{X}{(\log{}X)^{\alpha}}\right).$
@@ 2.2. Radu’s Algorithm
indexed by all the positive divisors $\delta$ of $M$, where $1=\delta_{1}<\delta_{2}<\cdots<\delta_{k}=M$. For positive integers $m$, and any integer $s$, let $[s]_{m}$ denote the residue class of $s$ in $\mathbb{Z}_{m}:=\mathbb{Z}/{m\mathbb{Z}}$. Let $\mathbb{Z}_{m}^{*}$ be the set of all invertible elements in $\mathbb{Z}_{m}$. Let $\mathbb{S}_{m}\subseteq\mathbb{Z}_{m}$ be the set of all squares in $\mathbb{Z}_{m}^{*}$. For $(r_{\delta})\in R(M)$ and $t\in\{0,1,\ldots,m-1\}$, we define a subset $P(t)\subseteq\{0,1,\ldots,m-1\}$ by
(2.2) $P(t):=\left\{t^{\prime}\in\{0,1,\ldots,m-1\}:t^{\prime}\equiv ts+\frac{s-1}{24}\sum_{\delta|M}\delta r_{\delta}\pmod{m},\text{ for some }[s]_{24m}\in\mathbb{S}_{24m}\right\}.$
@@
    If $2|m$, then either $4|kN$ and $8|sN$ or $2|s$ and $8|(1-j)N$, where $\prod_{\delta|M}\delta^{|r_{\delta}|}=2^{s}\cdot j$, here $s$ and $j$ are nonnegative integers with $j$ odd.
@@
Using Theorem 2.2, we see that $G_{1}(z)$ is a modular form of weight $390$, and level $4$, so $G_{1}(z)\in\mathcal{M}_{390}(\Gamma_{0}(4))$. From the definition of $a_{c}(n)$, we have
@@
We know that $G_{1}(z)\mid T_{43}\in\mathcal{M}_{390}(\Gamma_{0}(4))$, By Theorem 2.3, the Sturm bound for this space of forms is $195$. Using Mathematica, we verify that the coefficients of $q$-expansion of $G_{1}(z)\mid T_{43}$ up to the desired bound are congruent to 00 $\pmod{43}$. Thus, $G_{1}(z)\mid T_{43}\equiv 0\pmod{43}$, which proves (1.3).
@@
is Ramanujan’s $\varphi$-function. We can then iterate (4.1) to prove the next theorem.
@@
Using Lemma 4.1, we now prove Theorems 1.4 and 1.6.
@@
From the Binomial theorem, it is clear that $\varphi(q^{2i})^{t\cdot 2^{i-1}}\equiv 1\pmod{4}$ for each $i\geq 2$. Thus,
@@
Proof of Theorem 1.6.
From the Binomial theorem, it is clear that $\varphi(q^{2i})^{t\cdot 2^{i-1}}\equiv 1\pmod{8}$ for each $i\geq 3$. Thus,
$\equiv\varphi(q)\varphi(q^{2})^{c+1}\varphi(q^{4})^{2(c+1)}\pmod{8}$
$=\left(1+2\sum_{n\geq 1}q^{n^{2}}\right)\left(1+2\sum_{n\geq 1}q^{2n^{2}}\right)^{c+1}\left(1+2\sum_{n\geq 1}q^{4n^{2}}\right)^{2(c+1)}$
$=\left(1+2\sum_{n\geq 1}q^{n^{2}}\right)\left(\sum_{k=0}^{c+1}\binom{c+1}{k}2^{k}\left(\sum_{n\geq 1}q^{2n^{2}}\right)^{k}\right)\left(\sum_{k=0}^{2(c+1)}\binom{2(c+1)}{k}2^{k}\left(\sum_{n\geq 1}q^{4n^{2}}\right)^{k}\right)$
$\equiv\left(1+2\sum_{n\geq 1}q^{n^{2}}\right)\left(1+2(c+1)\sum_{n\geq 1}q^{2n^{2}}+2c(c+1)\sum_{n\geq 1}q^{4n^{2}}\right)$
${\hskip 12.0pt}\left(1+4(c+1)\sum_{n\geq 1}q^{4n^{2}}+4(c+1)\sum_{n\geq 1}q^{8n^{2}}\right)\pmod{8}$
$\equiv 1+2\sum_{n\geq 1}q^{n^{2}}+2(c+1)\sum_{n\geq 1}q^{2(2n-1)^{2}}+6(c+1)\sum_{n\geq 1}q^{2(2n)^{2}}$
${\hskip 12.0pt}+2(c+1)(c+2)\sum_{n\geq 1}q^{4n^{2}}+4(c+1)\sum_{m,n\geq 1}q^{n^{2}+2m^{2}}\pmod{8}.$
Hence, the result follows from the above identity. ∎
5. Proofs of Theorems 1.8, 1.9, 1.10, 1.11, and 1.12
@@
(5.1) $\varphi(q)$ $=\varphi\left(q^{4}\right)+2q\psi\left(q^{8}\right),$
(5.2) $f_{1}^{2}$ $=\frac{f_{2}f_{8}^{5}}{f_{4}^{2}f_{16}^{2}}-2q\frac{f_{2}f_{16}^{2}}{f_{8}},$
(5.3) $\frac{1}{f_{1}^{2}}$ $=\frac{f_{8}^{5}}{f_{2}^{5}f_{16}^{2}}+2q\frac{f_{4}^{2}f_{16}^{2}}{f_{2}^{5}f_{8}},$
(5.4) $\varphi(q)$ $=\varphi\left(q^{9}\right)+2q\Omega\left(q^{3}\right),$
(5.5) $\dfrac{f_{2}^{2}}{f_{1}}$ $=\frac{f_{6}f_{9}^{2}}{f_{3}f_{18}}+q\frac{f_{18}^{2}}{f_{9}},$
(5.6) $\dfrac{f_{1}}{f_{4}}$ $=\frac{f_{6}f_{9}f_{18}}{f_{12}^{3}}-q\frac{f_{3}f_{18}^{4}}{f_{9}^{2}f_{12}^{3}}-q^{2}\frac{f_{6}^{2}f_{9}f_{36}^{3}}{f_{12}^{4}f_{18}^{2}},$
@@
(5.1) and (5.3) are the same identity in [Hir17, Eq. (1.9.4)], (5.2) can be obtained by replacing $q$ by $-q$ in (5.1), while (5.4) and (5.5) are [Hir17, Eq. (26.1.1), (26.1.2)]. We prove (5.6) below. We need the following identities
(5.7) $\frac{f_{6}f_{12}^{2}f_{18}^{2}f_{36}^{2}}{f_{3}}$ $=\frac{f_{18}^{9}}{f_{9}^{3}}+q^{3}\frac{f_{6}^{3}f_{36}^{6}}{f_{3}f_{12}^{2}},$
(5.8) $f_{9}^{3}f_{12}$ $=\frac{f_{3}f_{12}^{4}f_{18}^{2}}{f_{6}^{2}f_{36}}+q^{3}f_{3}f_{36}^{3},$
(5.9) $\frac{f_{1}}{f_{2}^{2}}$ $=\frac{f_{3}^{2}f_{9}^{3}}{f_{6}^{6}}-q\frac{f_{3}^{3}f_{18}^{3}}{f_{6}^{7}}+q^{2}\frac{f_{3}^{4}f_{18}^{6}}{f_{6}^{8}f_{9}^{3}},$
(5.10) $\frac{f_{1}f_{4}}{f_{2}}$ $=\frac{f_{3}f_{12}f_{18}^{5}}{f_{6}^{2}f_{9}^{2}f_{36}^{2}}-q\frac{f_{9}f_{36}}{f_{18}}.$
(5.8) and (5.9) are [Hir17, Eqs. (22.7.5) and (37.2.3)] in disguise, respectively. (5.7) and (5.10) can be obtained by replacing $q$ by $-q$ in (5.8) and (5.5), respectively. With the help of (5.9) and (5.10), we have
@@
Now, we are ready to prove Theorem 1.8.
Proof of Theorem 1.8.
Using (5.3) and Lemma 5.2, we have
$\sum_{n\geq 0}\bar{a}_{c}(2n+1)q^{n}$ $=2\frac{f_{8}^{2}}{f_{4}}\left(\frac{f_{2}}{f_{1}^{2}}\right)^{c+1}=2\frac{f_{8}^{2}f_{2}^{c+1}}{f_{4}}\left(\frac{f_{8}^{5}}{f_{2}^{5}f_{16}^{2}}+2q\frac{f_{4}^{2}f_{16}^{2}}{f_{2}^{5}f_{8}}\right)^{c+1}$
$=2\frac{f_{8}^{2}f_{2}^{c+1}}{f_{4}}\sum_{k=0}^{c+1}\binom{c+1}{k}2^{k}q^{k}\left(\frac{f_{4}^{2}f_{16}^{2}}{f_{2}^{5}f_{8}}\right)^{k}\left(\frac{f_{8}^{5}}{F_{2}^{5}F_{16}^{2}}\right)^{c+1-k}$
(5.11) $=2\frac{f_{8}^{5c+7}}{f_{2}^{4c+4}f_{16}^{2c+2}f_{4}}\sum_{k=0}^{c+1}2^{k}q^{k}\binom{c+1}{k}\frac{f_{4}^{2k}f_{16}^{4k}}{f_{8}^{6k}}.$
From the above identity, extracting the terms with even and odd exponents of $q$ and then letting $q^{2}\rightarrow q$ and $c=2i$ for $i\geq 1$, we arrive at
(5.12) $\sum_{n\geq 0}\bar{a}_{2i}(4n+1)q^{n}$ $=2\frac{f_{4}^{10i+7}}{f_{1}^{8i+4}f_{2}f_{8}^{4i+2}}\sum_{k=0}^{i}\binom{2i+1}{2k}2^{2k}q^{k}\frac{f_{2}^{4k}f_{8}^{8k}}{f_{4}^{12k}},$
(5.13) $\sum_{n\geq 0}\bar{a}_{2i}(4n+3)q^{n}$ $=4\frac{f_{2}f_{4}^{10i+1}}{f_{1}^{8i+4}f_{8}^{4i-2}}\sum_{k=0}^{i}\binom{2i+1}{2k+1}2^{2k}q^{k}\frac{f_{2}^{4k}f_{8}^{8k}}{f_{4}^{12k}}.$
Equation (5.13) proves the congruence (1.11).
By a similar process, starting from the generating function for $\bar{a}_{c}(2n)$ from Lemma 5.2, we can arrive at
(5.14) $\sum_{n\geq 0}\bar{a}_{2i}(4n+2)q^{n}=2\frac{f_{2}^{7}f_{4}^{10i+1}}{f_{1}^{8i+6}f_{8}^{4i-2}}\sum_{k=0}^{i}2^{2k}q^{k}\binom{2i+1}{2k+1}\frac{f_{2}^{4k}f_{8}^{8k}}{f_{4}^{12k}}.$
Using (5.14) and (5.3), and then extracting the coefficients of odd powers of $q$, we obtain
$\sum_{n\geq 0}\bar{a}_{2i}(8n+6)q^{n}\equiv 0\pmod{4}.$
This clearly proves (1.12).
Similarly, using (5.12), (5.13) and (5.3), and then extracting the coefficients of odd powers of $q$, by a similar process, we prove that
$\sum_{n\geq 0}\bar{a}_{2i}(8n+5)q^{n}\equiv 0\pmod{4},$
$\sum_{n\geq 0}\bar{a}_{2i}(8n+7)q^{n}\equiv 0\pmod{8}.$
These above prove (1.13) and (1.14). ∎
Proof of Theorem 1.9.
This follows easily by observing from (1.2) and Lemma 5.2 that
$\bar{a}_{2i}(2n)\equiv\bar{a}_{2i}(n)\pmod{4}.$
Together with (1.11), (1.12), the above gives the result. ∎
@@
(5.15) $\sum_{n\geq 0}\bar{a}_{2^{k}i-2^{k-1}-2}(4n+1)q^{n}$ $\equiv 2\sum_{r=0}^{2^{k-2}}2^{2r}\binom{2^{k-1}+1}{2r}q^{r}\dfrac{f_{4}^{5\times 2^{k-1}-12r+7}}{f_{2}^{2^{k}-4r+3}f_{8}^{2^{k}-8r+2}}\pmod{2^{k+2}},$
(5.16) $\sum_{n\geq 0}\bar{a}_{2^{k}i-2^{k-1}-2}(4n+3)q^{n}$ $\equiv-4\sum_{r=0}^{2^{k-2}}2^{2r}\binom{2^{k-1}+1}{2r+1}q^{r}\dfrac{f_{4}^{5\times 2^{k-1}-12r+1}}{f_{2}^{2^{k}-4r+1}f_{8}^{2^{k}-8r-2}}\pmod{2^{k+2}}.$
@@
(5.17) $\sum_{n\geq 0}\bar{a}_{2^{k}i+2^{k-1}-3}(2n+1)q^{n}$ $\equiv 2\dfrac{f_{2}^{2^{k}i+2^{k-1}-2}f_{8}^{2}}{f_{1}^{2^{k+1}i+2^{k}-4}f_{4}}=2\dfrac{\psi\left(q^{4}\right)}{\varphi\left(-q\right)^{2^{k}i+2^{k-1}-2}}\pmod{2^{k+3}}$
(5.18) $\equiv 2\dfrac{\varphi\left(-q\right)^{2^{k-1}(6i+7)+2}\psi\left(q^{4}\right)}{\varphi\left(-q\right)^{2^{k+2}(i+1)}}\equiv 2\dfrac{\varphi\left(-q\right)^{2^{k-1}(6i+7)+2}\psi\left(q^{4}\right)}{\varphi\left(-q^{2}\right)^{2^{k+1}(i+1)}}\pmod{2^{k+3}}.$
@@
(5.19) ×$\quad\times\varphi\left(q^{2}\right)^{2^{k-1}(6i+7)-2r+1}\psi\left(q^{4}\right)^{2r+1}\pmod{2^{k+3}}.$

### v2

@@ Abstract.
Corresponding Author: Manjil P. Saikia (manjil.saikia@ahduni.edu.in)
@@ 1. Introduction
$a_{37}(43n+12)$ $\equiv 0\pmod{43}$, (1.3)
$a_{41}(47n+21)$ $\equiv 0\pmod{47}$, (1.4)
$a_{53}(59n+56)$ $\equiv 0\pmod{59}$, (1.5)
$a_{61}(67n+19)$ $\equiv 0\pmod{67}$, (1.6)
$a_{65}(71n+32)$ $\equiv 0\pmod{71}$, (1.7)
$a_{73}(79n+62)$ $\equiv 0\pmod{79}$, (1.8)
$a_{77}(83n+79)$ $\equiv 0\pmod{83}$. (1.9)
@@
After we finished the proof of Theorem 1.1, Guadalupe [Gua25] updated his preprint on arXiv, which actually generalizes the above result. However, he uses elementary techniques to achieve this, so we keep our proof in this article.
@@
We also give a complete characterization of the $\bar{a}_{c}(n)$ function modulo $4$ in the following results.
@@
Theorem 1.4 is proved in Section 4 using elementary techniques.
@@
$\bar{a}_{2^{k}i-2^{k-1}-2}(8n+5)$ $\equiv 0\pmod{2^{k+1}}$, (1.11)
$\bar{a}_{2^{k}i-2^{k-1}-2}(8n+7)$ $\equiv 0\pmod{2^{k+2}}$. (1.12)
@@
$\bar{a}_{2^{k}i+\frac{2^{k}-6}{2}}(8n+7)$ $\equiv 0\pmod{2^{k+3}}$, (1.13)
$\bar{a}_{2^{k}i+\frac{2^{k}-2}{2}}(8n+7)$ $\equiv 0\pmod{2^{k+3}}$. (1.14)
@@
$\bar{a}_{9i}(9n+6)$ $\equiv 0\pmod{12}$, (1.15)
$\bar{a}_{9i+4}(9n+6)$ $\equiv 0\pmod{12}$, (1.16)
$\bar{a}_{9i+8}(9n+6)$ $\equiv 0\pmod{12}$. (1.17)
Theorems 1.6, 1.7, and 1.8 are proved in Section 5 using elementary techniques.
@@
which means that almost all the coefficients of $A$ are divisible by $M$. It turns out that the partition function $\bar{a}_{c}(n)$ also satisfies such results. Specifically, we prove the following results in this article.
@@
$\{n\in\mathbb{N}:\bar{a}_{2^{\alpha}m}(n)\equiv 0\pmod{2^{k}}\}$
@@
$\lim_{X\to\infty}\frac{\#\{0\leq n\leq X:\bar{a}_{2^{\alpha}m}(n)\equiv 0\pmod{2^{k}}\}}{X}=1.$
@@
$\{n\in\mathbb{N}:\bar{a}_{2^{\alpha}m}(n)\equiv 0\pmod{3^{k}}\}$
@@
$\lim_{X\to\infty}\frac{\#\{0\leq n\leq X:\bar{a}_{2^{\alpha}m}(n)\equiv 0\pmod{3^{k}}\}}{X}=1.$
@@
Theorems 1.9, 1.10, and 1.11 are proved in Section 6.
@@ 2.1. Various results on modular forms
$\Gamma$ $:=\left\{\begin{bmatrix}a&b\\ c&d\end{bmatrix}:a,b,c,d\in\mathbb{Z},ad-bc=1\right\},$
$\Gamma_{\infty}$ $:=\left\{\begin{bmatrix}1&n\\ 0&1\end{bmatrix}\in\Gamma:n\in\mathbb{Z}\right\}.$
@@
$\Gamma_{0}(N):=\left\{\begin{bmatrix}a&b\\ c&d\end{bmatrix}\in\Gamma:c\equiv~0\pmod{N}\right\}$,
$\Gamma_{1}(N):=\left\{\begin{bmatrix}a&b\\ c&d\end{bmatrix}\in\Gamma_{0}(N):a\equiv d\equiv~1\pmod{N}\right\}$.
@@
If $\chi$ is a Dirichlet character modulo $N$, then a form $f(z)\in M_{l}(\Gamma_{1}(N))$ has Nebentypus character $\chi$ if
$f\left(\frac{az+b}{cz+d}\right)=\chi(d)(cz+d)^{\ell}f(z),$
@@
Let $k$ be an integer and $g(z)=\sum_{n=0}^{\infty}a(n)q^{n}$ a modular form of weight $k$ for $\Gamma_{0}(N)$. For any given positive integer $m$, if $a(n)\equiv 0\pmod{m}$ holds for all $n\leq\dfrac{kN}{12}\prod\limits_{p~\text{prime},~p\,|\,N}~\left(1+\dfrac{1}{p}\right)\,$ then $a(n)\equiv 0\pmod{m}$ holds for any $n\geq 0$.
@@
$f(z)\,|\,T_{p}:=\sum_{n=0}^{\infty}\left(a(pn)+\chi(p)p^{\ell-1}a\left(\frac{n}{p}\right)\right)q^{n}.$ (2.1)
@@
$\left\{0<n\leq X:b(n)\not\equiv 0\pmod{r}\right\}=\mathcal{O}\left(\frac{X}{(\log{}X)^{\alpha}}\right).$
@@ 2.2. Radu’s Algorithm
indexed by all the positive divisors $\delta$ of $M$, where $1=\delta_{1}<\delta_{2}<\cdots<\delta_{k}=M$. For positive integers $m$, and any integer $s$, let $s]_{m}$ denote the residue class of $s$ in $\mathbb{Z}_{m}:=\mathbb{Z}/{m\mathbb{Z}}$. Let $\mathbb{Z}_{m}^{*}$ be the set of all invertible elements in $\mathbb{Z}_{m}$. Let $\mathbb{S}_{m}\subseteq\mathbb{Z}_{m}$ be the set of all squares in $\mathbb{Z}_{m}^{*}$. For $(r_{\delta})\in R(M)$ and $t\in\{0,1,\ldots,m-1\}$, we define a subset $P(t)\subseteq\{0,1,\ldots,m-1\}$ by
[…]
 $\quad\left.\text{ for some }[s]_{24m}\in\mathbb{S}_{24m}\right\}.$ (2.2)
@@
    If $2|m$, then either $4|kN$ and $8|sN$ or $2|s$ and 8|(1j)N8|(1-j)N, where $\prod_{\delta|M}\delta^{|r_{\delta}|}=2^{s}\cdot j$, here $s$ and $j$ are nonnegative integers with $j$ odd.
@@
Using Theorem 2.2, we see that $G_{1}(z)$ is a modular form of weight $390$, and level $4$, so $G_{1}(z)\in{M}_{390}(\Gamma_{0}(4))$. From the definition of $a_{c}(n)$, we have
@@
We know that $G_{1}(z)\mid T_{43}\in{M}_{390}(\Gamma_{0}(4))$, By Theorem 2.3, the Sturm bound for this space of forms is $195$. Using Mathematica, we verify that the coefficients of $q$-expansion of $G_{1}(z)\mid T_{43}$ up to the desired bound are congruent to 0 $\pmod{43}$. Thus, $G_{1}(z)\mid T_{43}\equiv 0\pmod{43}$, which proves (1.3).
@@
Congruences $\eta$-quotient Level Weight Sturm bound
(1.4) $\dfrac{\eta(z)^{704}}{\eta(2z)^{40}}$ 2 332 83
(1.5) $\dfrac{\eta(z)^{176}}{\eta(2z)^{52}}$ 4 62 31
(1.6) $\dfrac{\eta(z)^{1272}}{\eta(2z)^{60}}$ 4 606 303
(1.7) $\dfrac{\eta(z)^{1064}}{\eta(2z)^{64}}$ 2 500 125
(1.8) $\dfrac{\eta(z)^{552}}{\eta(2z)^{72}}$ 2 240 60
(1.9) $\dfrac{\eta(z)^{248}}{\eta(2z)^{76}}$ 4 86 43
@@
is Ramanujan’s $\varphi$-function. We can then iterate (4.1) to prove the next lemma.
@@
Using Lemma 4.1, we now prove Theorem 1.4.
@@
From the Binomial theorem, it is clear that $\varphi(q^{2^{i}})^{t\cdot 2^{i-1}}\equiv 1\pmod{4}$ for each $i\geq 2$. Thus,
@@
5. Proofs of Theorems 1.6, 1.7, and 1.8
@@
$\varphi(q)$ $=\varphi\left(q^{4}\right)+2q\psi\left(q^{8}\right),$ (5.1)
$f_{1}^{2}$ $=\frac{f_{2}f_{8}^{5}}{f_{4}^{2}f_{16}^{2}}-2q\frac{f_{2}f_{16}^{2}}{f_{8}},$ (5.2)
$\frac{1}{f_{1}^{2}}$ $=\frac{f_{8}^{5}}{f_{2}^{5}f_{16}^{2}}+2q\frac{f_{4}^{2}f_{16}^{2}}{f_{2}^{5}f_{8}},$ (5.3)
$\varphi(q)$ $=\varphi\left(q^{9}\right)+2q\Omega\left(q^{3}\right),$ (5.4)
$\dfrac{f_{2}^{2}}{f_{1}}$ $=\frac{f_{6}f_{9}^{2}}{f_{3}f_{18}}+q\frac{f_{18}^{2}}{f_{9}},$ (5.5)
$\dfrac{f_{1}}{f_{4}}$ $=\frac{f_{6}f_{9}f_{18}}{f_{12}^{3}}-q\frac{f_{3}f_{18}^{4}}{f_{9}^{2}f_{12}^{3}}-q^{2}\frac{f_{6}^{2}f_{9}f_{36}^{3}}{f_{12}^{4}f_{18}^{2}},$ (5.6)
@@
(5.1) and (5.3) are the same identity in [Hir17, Eq. (1.9.4)], (5.2) can be obtained by replacing $q$ by q-q in (5.1), while (5.4) and (5.5) are [Hir17, Eq. (26.1.1), (26.1.2)]. We prove (5.6) below. We need the following identities
$\frac{f_{6}f_{12}^{2}f_{18}^{2}f_{36}^{2}}{f_{3}}$ $=\frac{f_{18}^{9}}{f_{9}^{3}}+q^{3}\frac{f_{6}^{3}f_{36}^{6}}{f_{3}f_{12}^{2}},$ (5.7)
$f_{9}^{3}f_{12}$ $=\frac{f_{3}f_{12}^{4}f_{18}^{2}}{f_{6}^{2}f_{36}}+q^{3}f_{3}f_{36}^{3},$ (5.8)
$\frac{f_{1}}{f_{2}^{2}}$ $=\frac{f_{3}^{2}f_{9}^{3}}{f_{6}^{6}}-q\frac{f_{3}^{3}f_{18}^{3}}{f_{6}^{7}}+q^{2}\frac{f_{3}^{4}f_{18}^{6}}{f_{6}^{8}f_{9}^{3}},$ (5.9)
$\frac{f_{1}f_{4}}{f_{2}}$ $=\frac{f_{3}f_{12}f_{18}^{5}}{f_{6}^{2}f_{9}^{2}f_{36}^{2}}-q\frac{f_{9}f_{36}}{f_{18}}.$ (5.10)
(5.8) and (5.9) are [Hir17, Eqs. (22.7.5) and (37.2.3)] in disguise, respectively. (5.7) and (5.10) can be obtained by replacing $q$ by q-q in (5.8) and (5.5), respectively. With the help of (5.9) and (5.10), we have
@@
$\sum_{n\geq 0}\bar{a}_{2^{k}i-2^{k-1}-2}(4n+1)q^{n}$ $\equiv 2\sum_{r=0}^{2^{k-2}}2^{2r}\binom{2^{k-1}+1}{2r}q^{r}\dfrac{f_{4}^{5\times 2^{k-1}-12r+7}}{f_{2}^{2^{k}-4r+3}f_{8}^{2^{k}-8r+2}}\pmod{2^{k+2}}$, (5.11)
$\sum_{n\geq 0}\bar{a}_{2^{k}i-2^{k-1}-2}(4n+3)q^{n}$ $\equiv-4\sum_{r=0}^{2^{k-2}}2^{2r}\binom{2^{k-1}+1}{2r+1}q^{r}\dfrac{f_{4}^{5\times 2^{k-1}-12r+1}}{f_{2}^{2^{k}-4r+1}f_{8}^{2^{k}-8r-2}}\pmod{2^{k+2}}$. (5.12)
@@
$\sum_{n\geq 0}\bar{a}_{2^{k}i+2^{k-1}-3}(2n+1)q^{n}$ $\equiv 2\dfrac{f_{2}^{2^{k}i+2^{k-1}-2}f_{8}^{2}}{f_{1}^{2^{k+1}i+2^{k}-4}f_{4}}=2\dfrac{\psi\left(q^{4}\right)}{\varphi\left(-q\right)^{2^{k}i+2^{k-1}-2}}\pmod{2^{k+3}}$ (5.13)
$\equiv 2\dfrac{\varphi\left(-q\right)^{2^{k-1}(6i+7)+2}\psi\left(q^{4}\right)}{\varphi\left(-q\right)^{2^{k+2}(i+1)}}$
$\equiv 2\dfrac{\varphi\left(-q\right)^{2^{k-1}(6i+7)+2}\psi\left(q^{4}\right)}{\varphi\left(-q^{2}\right)^{2^{k+1}(i+1)}}\pmod{2^{k+3}}$. (5.14)
@@
$\quad\times\varphi\left(q^{2}\right)^{2^{k-1}(6i+7)-2r+1}\psi\left(q^{4}\right)^{2r+1}\pmod{2^{k+3}}$. (5.15)